OpenStudy (anonymous):

What is the center of the circle with equation (x + 2)² + (y - 5)² = 16

OpenStudy (primeralph):

-2,5

OpenStudy (nurali):

Standard equation of circle: (x - h)^2 + (y - k)^2 = r^2, where (h, k) is the centre point and r is the radius Rewrite the equation in standard form: (x + 2)² + (y - 5)² = 16 Now you just match it up with the standard equation values. h = -2, k = 5, and r^2 = 16 Therefore the centre is point (-2, 5) and the radius is √16 = 4.

OpenStudy (primeralph):

I tried to come back to explain, but the stuff kept logging me out.

OpenStudy (primeralph):